Feb 12, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items


Tentative Agenda

  1. Announcements
  2. Collaboration Meeting February 19-21 at Jefferson Lab.
  3. Action Items
  4. FCAL
    1. HDFCAL log book
    2. Calibration Update
  5. BCAL HDBCAL log book
  6. Reconstruction and Simulation
  7. Calibration
  8. Any other business


Attendees: Elton, Will, Noemi, Adesh, Manuel, Mark, Eugene (JLab); Zisis (UofR); Matt (IU); Christina, George (Athens)

  1. Announcements
  2. Collaboration Meeting February 19-21 at Jefferson Lab.
  3. Action Items
  4. FCAL
    1. HDFCAL log book
    2. Hardware updates (Manuel)
      • GUIs for controls will be completed this week. These incorporates changes discussed with Hovanes for good interface to EPICS. The EPICS GUIs will allow entire strands to be turned on at once instead of one pmt at a time.
      • Testing of control software will be done next week.
      • Dark room environment was tested after all the VESDA work. No new light leaks have been found.
      • Replaced about 20 bases. Shipped back to IU for diagnosis. Do not expect to replace any more bases before next run.
      • New checklists need to be completed starting next week in preparation for the run.
    3. Calibration Update (Matt)
      • Emphasis on updates since yesterday's calibration meeting.
      • Gain matching by minimizing the pi0 width. Iterations have improved width but also distort the distribution by gathering in background into the peak region.
      • Checking systematics by dividing the sample into one: one for determining the gain constants and the other for testing their validity
      • Looking at alternate methods for equalizing the gains.
      • Response to LED. Set gains based on average of nearest neighbors. Variations on this scheme have been tried and testing vs number of iterations, were a large number of iteration is not optimal.
      • First outcome of LED smoothing algorithms seems to narrow the pi0 peak, but not as much as the pi0 minimization procedure.
      • Investigating the use of electrons (E/p) method to determine constants.
    4. Comments (Eugene)
      • Talked to software reviewer from GAMS that is calibration 3500 LGD array. They have an iterative method of plotting the pi0 peak with one photon in a particular block and use this to adjust the gain for that block.
      • They determine the calibration constants in 5 energy ranges (periphery) and 20 energy ranges (center of cal) to determine non-linearities.
      • [Note added] Eugene: estimates the pi0 mass width should be about 10 MeV at 2 GeV.
      • Christina: Do you have sufficient statistics to investigate the pi0 peak in different regions of the calorimeter? Matt: No, but all photons are selected so that they are at least a couple of blocks away from the periphery (beam hole or outer radius).
  5. BCAL HDBCAL log book
    • BCAL is operational. Elton and Noemi are working on the pulser system to check status and complete the checklist.
  6. Reconstruction and Simulation
  7. Andrei and Tegan are working on understanding the "curvature" effects of showers in the bcal.
  8. Calibration (Elton)
    1. Voltage settings for Down 14, 15 and 16. See Log Entry 3319846
      • For these modules, the voltage was ofter set incorrectly to 75 volts, instead of the nominal 1.2 V above bias. Calibrations will need to make allowances for the changes as a function of time. The periods are identified in the log entries.
    2. Gain calibration constants Up*Down and Up/Down (Elton)
      • The pi0 width is mainly sensitive only to the combination U*D. The ratio U/D is sensitive to the position of the shower determined from time measurements
      • In the next iteration, Will's program will only be used to extract U*D and Mark's program to measure attenuation lengths will determine U/D.
    3. Thoughts on veff and time-walk constants
      • George and Christina will be looking into the determination of veff and time walk constants starting in mid March and clarify any issues during George's 2 week stay at JLab in mid April. Work can continue once he returns to Athens.
    4. Pi0 width minimization (Will)
      • A complete iteration of the minimization scheme has been completed. Seems to narrow the width by about 10%. Sample used had Eg>0.75 GeV with pi0/Back~4/1 purity.
      • Currently working on obtaining a skim enriched in pi0s to facilitate and speed up any iterations with the width minimization.
    5. Attenuation lengths (Mark)
      • Extracting attenuation lengths for each file from commissioning.
      • Finds some systematic changes in the attenuation lengths mainly in mode 7 running.
      • Unclear why there is a difference between the lambda determination in mode 7 vs mode 8. Differences are of order delta_lambda/lambda ~ 10%.
      • Investigations are ongoing to understand the source.
  9. Any other business